Graph each pair of equations on a graphing calculator using a window that shows the point of intersection. Use the intesect feature to find the solution
3x - 2y = -158
5x + 4y = 1526

Answer by cartert(27)

first you need to simplify the equations
(A) 3x - 2y = -158 (B) 5x + 4y = 1526
-3x -3x -5x -5x
-2y = -3x - 158 4y = -5x + 1526
-2y/-2 = -3x/-2 - 158/-2 4y/4 = -5x/4 + 1526/4
Answer-- y = 3/2x + 79 Answer-- y = -5/4x + 381 1/2
then graph
the points for (A) are {(-30,34),(-10,64),(10,94)}
the points for (B) are {(282,29),(6,374),(-198,629)}
the intersection point is (110,244)
I really hope this helps thank you for using www.algebra.com